# SCARIF: Towards Carbon Modeling of Cloud Servers with Accelerators

### v2

@@ IV-C Embodied Carbon Cost Modeling for Accelerators
Hardware accelerators such as GPUs and FPGAs are important in today’s computing systems, particularly for servers in modern datacenters. However, none of the vendors provide the carbon footprint report data for servers equipped with accelerators, making it hard to analyze the carbon cost of the computing systems equipped with the accelerators. Unfortunately, GPU and FPGA vendors are not as forthcoming with sustainability data. Here, we propose a method to roughly estimate the carbon cost of such systems by utilizing the ACT tool and some scaling assumptions that can address missing costs that are proportional to the complexity of the accelerator.
Different from CPUs, there are fewer models of accelerators but more varieties of architectures among the models. As a result, a unified feature like CPU core numbers is impractical for accelerators since the GPU core usually takes a different area from the CPU core, and there is no real core equivalent in an FPGA. Here, we assume that the GPU-related part, which contains not only the chip itself but also the peripheral components like the slots, shares the same pattern as the CPU-related part. Thus, the carbon of the GPU-related part can be obtained via the carbon of the CPU-related part and the chip area of both parts. We leverage the ACT tool in generating the estimation. When the model of the CPU and GPU are specified, the carbon cost of their chip computed by ACT, and can be obtained. Also, using the generated SCARIF model, we can get the carbon cost of CPU-related part as . Thus the GPU part carbon cost can be roughly estimated as:
@@ V Carbon Analysis From SCARIF
In this section, we show how can we obtain a holistic view of the server carbon footprints by using SCARIF for several different scenarios under given workloads. We provide concrete case studies and analysis: (1) to compare the total carbon costs, including both manufacturing and operational costs, for systems with or without GPUs; (2) to demonstrate the breakeven analysis when upgrading servers with newer generation GPUs; (3) to compare system when choosing different kinds of accelerators, i.e., GPUs or FPGAs.
@@ V-A Analysis and Experimental Setup
Hardware configurations. We use two generations of servers in our analysis, Dell R740 and Dell R750, released in 2017 and 2020 respectively. For Dell R740, we configure 2x Intel 14nm Xeon 8180 CPUs with 694 chip area, 64GB DDR4-10nm DRAM, and 1TB HDD. For Dell R750, we configure 2x Intel 10nm Xeon 8375 CPUs with 660 chip area, 64GB DDR4-10nm DRAM, and 1TB HDD. Here the configurations are the two real system configurations that we use for real-world application workload performance and energy characterization. For the accelerators, we choose NVIDIA V100 GPU and NVIDIA A100 GPU, the 12nm V100 was released in 2017, with a chip area of 815 , whereas the 7nm A100 was released in 2020, with a chip area of 826 . We also choose the Xilinx ZCU102 FPGA for comparison, which was released in 2016, has a fabrication of 16nm, and a chip area of 245 .

### v3

@@ IV-C Embodied Carbon Cost Modeling for Accelerators
Hardware accelerators such as GPUs and FPGAs are important in today’s computing systems, particularly in modern datacenters. However, none of the vendors provide the carbon footprint report data for servers equipped with accelerators, and the GPU and FPGA vendors are not as forthcoming with sustainability data. Here, we propose a method to roughly estimate the carbon cost of such systems by utilizing the ACT tool and some scaling assumptions of accelerators.
Different from CPUs, there are fewer models of accelerators but more varieties of architectures among the models. As a result, a unified feature like CPU core numbers is impractical for accelerators since the GPU core usually takes a different area from the CPU core, and there is no real core equivalent in an FPGA. Here, we assume that the GPU-related part, which contains not only the chip itself but also the peripheral components like the slots, shares the same pattern as the accelerator-related part. Thus, the carbon of the accelerator-related part can be obtained by analyzing the carbon of the CPU-related part and the chip area of both parts. We leverage the ACT tool in generating the estimation. Given the model of the accelerator, its carbon cost of the chip, can be obtained by ACT. Then the accelerator part carbon cost can be roughly estimated as:
Where is a hyper-parameter which can be obtained by integrating different CPU-only server setups as the baseline.
@@ V Carbon Analysis From SCARIF
In this section, we show how can we holisticly view the server carbon footprints by using SCARIF for several different scenarios under given workloads. We provide concrete case studies and analysis: (1) to compare both manufacturing and operational costs for systems; (2) to demonstrate the breakeven analysis when upgrading servers; (3) to compare system with different kinds of accelerators, i.e., GPUs or FPGAs.
@@ V-A Analysis and Experimental Setup
Hardware configurations. We use two generations of servers in our analysis, Dell R740 and Dell R750, released in 2017 and 2020 respectively. For Dell R740, we configure 2x Intel 14nm Xeon 8180 CPUs with 694 chip area, 64GB DDR4-10nm DRAM, and 1TB HDD. For Dell R750, we configure 2x Intel 10nm Xeon 8375 CPUs with 660 chip area, 64GB DDR4-10nm DRAM, and 1TB HDD. For the accelerators, we choose NVIDIA V100 GPU and NVIDIA A100 GPU, the 12nm V100 was released in 2017, with a chip area of 815 , whereas the 7nm A100 was released in 2020, with a chip area of 826 . We also choose the Xilinx ZCU102 FPGA for comparison, which was released in 2016, has a fabrication of 16nm, and a chip area of 245 .
@@ V-C Carbon Cost for Systems with GPU and FPGA
VI Conclusion and Acknowledge
SCARIF provides a key tool for modeling the embodied carbon cost of server systems with accelerators. We will add SCARIF with more features to further guide the system carbon emission optimization from edge-level to enterprise-level.
We acknowledge NSF support (#2213701, #2217003, #2324864, #2328972) and the University of Pittsburgh New Faculty Start-up Grant. We thank the AMD/Xilinx University Program for hardware and software donations.